# Synthetic Homes: An Accessible Multimodal Pipeline for Producing Residential Building Data with Generative AI

### v2

@@ 5 Conclusion
In this work, we presented a novel way to generate multimodal data for urban energy research and other fields that use similar data. Using a multimodal pipeline, we produce this data in a cost-effective and efficient way. We also performed an in-depth analysis to ensure that our generative AIpipeline components were behaving as expected, taking into account the concerns of [24].
Generative AIis a powerful tool, and we recognize the importance of this new frontier as described by Xu et al. [17] and Liu et al [24]. Not only does our work demonstrate a step forward in the urban energy field, it also presents a useful, malleable tool that experts and researchers can use for a variety of tasks. There is much work that can utilize or expand upon this study. One promising direction is the use of this pipeline to train MLalgorithms.
For instance, a new model of the pipeline could utilize the resulting simulation data and inspection notes to label the efficiency of particular components that affect the energy efficiency of a home. These labeled homes could be used as inputs for an MLmodel that could provide recommendations for energy efficiency retrofits. Furthermore, this model, trained with our synthetic data, could be used to predict the priorities for retrofits in real houses.
@@
Acknowledgments
The authors extend gratitude to Lafayette College for supporting this work by funding it through the EXCEL Scholars program and providing computational resources with the Firebird high performance computing cluster. Access to the JetStream2 system was made possible via the NSF-funded ACCESS program, facilitated by Lafayette College. We also thank the NSF and ACCESS for supporting this resource, and Indiana University for hosting JetStream2 as a part of the ACCESS program.
CRediT Author Attributions
Jackson Eshbaugh: Software, Investigation, Writing - Original Draft, Visualization, Validation, Data Curation. Chetan Tiwari: Conceptualization, Writing - Review and Editing, Supervision (supporting). Jorge Silveyra: Conceptualization, Writing - Review and Editing, Methodology, Supervision (lead), Project Administration, Validation
Glossary
@@ Appendix A Occlusion Results
Next, we provide our hypotheses pertaining to the performance of both models. We hypothesize that LLaVA’s advantage stems from its architecture: it employs an early fusion design in which images are converted into visual tokens by CLIP [46] that are then prepended to the textual prompt [28]. These visual tokens are passed through the transformer layers of the LLMtogether with text tokens, enabling multimodal attention at every layer. It has been demonstrated that vision tokens are most critical for a model in early layers, when they fuse into text tokens [47]. A likely result of this is the reduction of the model’s sensitivity to variations in the image, which may explain its stronger performance in our experiments.
While GPT-4V (GPT’s vision model) is proprietary, the case can be made that it possesses a different structure. Evaluation studies show that it performs well across visual reasoning and captioning tasks, but it falls short in fine grained visual problems, requiring complex visual tasks to be broken up into interleaved subfigures and text [48]. These evaluations are similar to evaluations from late fusion (connector) models. For example, BLIP-2 shows strengths in captioning images and instruction-following, while it is less robust with fine-grained spatial reasoning [49]. Another model, MiniGPT-4, demonstrates a similar profile: it has strengths in tasks such as image description, but falters with complex spatial reasoning [50]. Both BLIP-2 and MiniGPT-4 are late fusion models, where compressed visual features are routed into the LLMat later stages [49, 50]. As shown, BLIP-2, MiniGPT-4, and GPT-4V all suffer from a very similar weakness: fine-grained or complex spatial reasoning. Based on this similarity, we hypothesize that GPT-4V is also a late fusion model. This helps explain why GPTperformed differently from LLaVAin our experiments: as structure dictates function, differences in behavior strongly suggest differences in design. On the basis of this analysis, we feel confident to select LLaVAto complete this task.

### v3

@@ 5 Conclusion
In this work, we presented a novel way to generate multimodal building-scale data for urban energy research and other related fields. Using a multimodal pipeline, we produced this data in a cost-effective and efficient way. We also performed an in-depth analysis to ensure that our generative AIpipeline components were behaving as expected, taking into account the concerns of Liu et al. (2025).
Generative AIis a powerful tool, and we recognize the importance of this new frontier as described by Xu et al. (2024) and  Liu et al. (2025). Not only does our work demonstrate a step forward in the urban energy field, it also presents a useful, malleable tool that experts and researchers can use for a variety of tasks.
One promising direction is the use of this pipeline to train MLalgorithms. A natural expansion of the pipeline could utilize the resulting EnergyPlus simulation data and inspection notes to rank the efficiency of particular components that affect the energy efficiency of a home. These labeled homes could be used as inputs for an MLmodel that could provide recommendations for energy efficiency retrofits. Furthermore, this model, trained with our synthetic data, could be used to predict the priorities for retrofits in real houses.
@@ Appendix A Occlusion Results
Next, we provide our hypotheses pertaining to the performance of both models. We hypothesize that LLaVA’s advantage stems from its architecture: it employs an early fusion design in which images are converted into visual tokens by CLIP (Radford et al., 2021) that are then prepended to the textual prompt (Liu et al., 2023). These visual tokens are passed through the transformer layers of the LLMtogether with text tokens, enabling multimodal attention at every layer. It has been demonstrated that vision tokens are most critical for a model in early layers, when they fuse into text tokens (Zhang et al., 2025b). A likely result of this is the reduction of the model’s sensitivity to variations in the image, which may explain its stronger performance in our experiments.
While GPT-4V (GPT’s vision model) is proprietary, the case can be made that it possesses a different structure. Evaluation studies show that it performs well across visual reasoning and captioning tasks, but it falls short in fine grained visual problems, requiring complex visual tasks to be broken up into interleaved subfigures and text (Yang et al., 2023). These evaluations are similar to evaluations from late fusion (connector) models. For example, BLIP-2 shows strengths in captioning images and instruction-following, while it is less robust with fine-grained spatial reasoning (Li et al., 2023). Another model, MiniGPT-4, demonstrates a similar profile: it has strengths in tasks such as image description, but falters with complex spatial reasoning (Zhu et al., 2023). Both BLIP-2 and MiniGPT-4 are late fusion models, where compressed visual features are routed into the LLMat later stages (Li et al., 2023; Zhu et al., 2023). As shown, BLIP-2, MiniGPT-4, and GPT-4V all suffer from a very similar weakness: fine-grained or complex spatial reasoning. Based on this similarity, we hypothesize that GPT-4V is also a late fusion model. This helps explain why GPTperformed differently from LLaVAin our experiments: as structure dictates function, differences in behavior strongly suggest differences in design. On the basis of this analysis, we feel confident to select LLaVAto complete this task.